The height of a ball thrown upward from the top of a building is given by: h(t)=-16t^2+64t+80
What is the height of the ball at 1 second?
What is the maximum height reached by the ball?
How long will it take for the ball to reach the ground?

Answer by stanbon(75887)

The height of a ball thrown upward from the top of a building is given by: h(t)=-16t^2+64t+80
What is the height of the ball at 1 second?
h(1) = -16+64+80 = 128 ft.
---------------------------------
What is the maximum height reached by the ball?
Max occurs when x = -b/(2a) = -64/(2*-16) = 2 sec
h(2) = -16(2^2) + 64(2) + 80 = 144 ft
----------------------
How long will it take for the ball to reach the ground?
When it reaches the ground the height is zero.
-------
Solve:: -16t^2+64t+80 = 0
t^2 - 4t - 5 = 0
(t-5)(t+1) = 0
Positive solution::
t = 5 seconds
============
Cheers,
Stan H.
----------------